# Beyond Pixels: Efficient Dataset Distillation via Sparse Gaussian Representation

### v1

@@ 4.1 Experimental Results
Storage Budget and Evaluation Protocol
We evaluate performance under varying storage budgets, specified in IPC (images per class). Since our method often utilizes multiple low-storage Gaussian images, we additionally report GPC (Gaussian Images per Class) in Appendix A.1. The number of Gaussians per image is computed as $\text{pts}=\frac{\text{res}\times\text{res}\times 3\times\text{IPC}\times 2}{\text{GPC}\times 9}$, where the factor of 2 accounts for our use of bf16 to store each Gaussian parameter, and the denominator 9 reflects the total number of parameters per Gaussian.
Performance Comparison
Through the integration of GSDD, our method consistently outperforms the pixel-based TM baseline and other dataset parameterization methods on six ImageNet subsets with IPC=1, as shown in Table 2. Extensive results on CIFAR-10 and CIFAR-100 with IPC values of 1, 10, and 50 are provided in Appendix A.2. When compared to the state-of-the-art method DDiF, our approach achieves consistent improvements, demonstrating highly competitive performance. These results highlight the effectiveness of GSDD in enhancing the quality of distilled datasets.
Table 4: Cross Architecture Results
Subset Nette Woof Fruit Yellow Meow Squawk
TM 22.0 14.8 17.1 22.3 16.2 25.5
IDC 27.9 19.5 23.9 28.0 19.8 29.9
FreD 36.2 23.7 23.6 31.2 19.1 37.4
GLaD 30.4 17.1 21.1 19.6 28.2
H-GLaD 30.8 17.4 21.5 20.1 28.8
LD3M 32.0 19.9 21.4 22.1 30.4
DDiF 59.3 34.1 39.3 51.1 33.8 54.0
GSDD 58.1 34.6 39.9 53.6 34.0 58.0
Universality to Matching Objectives
Cross-Architecture Performance
An essential property of dataset distillation is robustness across model architectures. The distilled data should retain high performance even when the evaluation model differs from the one used during distillation. To assess this, we synthesize data using a ConvNet and evaluate it on a variety of downstream architectures, including ResNet, VGG, and others. The results, summarized in Table 4, demonstrate that the increased diversity introduced by GSDD leads to substantial improvements in cross-architecture generalization.
4.2 More Analysis
@@
To investigate the contribution of each component to model performance, we conduct a series of ablation studies, with results summarized in Table 3. The experiments show that opacity modeling and bf16 mixed-precision training are two major factors, as removing either leads to a notable performance drop. The impact of anti-aliasing exhibits a dependence on the dataset and the complexity of the representation. Its effect becomes important when the number of Gaussians used to represent each image is limited. Specifically, for datasets like CIFAR-10, where storage budgets are constrained and each image is represented by a small number of Gaussians, disabling anti-aliasing introduces rendering artifacts that degrade performance. In contrast, for datasets like ImageNette with denser Gaussians, aliasing artifacts are largely mitigated, and the performance gains from anti-aliasing become relatively minor.
To comprehensively evaluate the efficiency of our approach, we conduct a quantitative comparison with the state-of-the-art DDiF method in terms of both decoding and training performance. As shown in Figure 5, we measure the inference time and memory consumption of both methods across varying per-image storage budgets, image resolutions, and batch sizes. The results demonstrate that GSDD consistently outperforms DDiF in both inference/training speed and memory efficiency. This advantage becomes especially pronounced when handling high-resolution inputs or large batch sizes, where GSDD achieves several-fold, or even an order-of-magnitude, reductions in both computation time and memory usage. These findings indicate that our method not only maintains strong image representation capabilities for distillation but also offers superior rendering speed and significantly lower memory footprint, thereby substantially improving the scalability of dataset distillation methods in practical deployment scenarios.
@@
Appendix A Appendix
A.1 Experimental Details
A.1.1 Datasets
@@ A.1.4 Dataset Distillation
Table 5: Hyperparameters for dataset distillation on CIFAR-10.
@@
Table 6: Hyperparameters for dataset distillation on CIFAR-100.
@@
Table 7: Hyperparameters for dataset distillation on ImageNet-subset TM).
@@
Table 8: Hyperparameters for dataset distillation on ImageNet-subset (DM).
@@
Table 9: Hyperparameters for dataset distillation on ImageNet-subset (DC).
@@
A.2 More Experimental Results
@@
We evaluate the generalization ability of distilled data across different architectures, including AlexNet, VGG11, ResNet18, and ViT. The network implementations follow those provided in Cazenavette et al. (2023). For training, the learning rate is set to $1\times 10^{-3}$ for AlexNet, $5\times 10^{-5}$ for ViT, and $1\times 10^{-2}$ for the remaining architectures. All models are optimized using Adam with a cosine annealing learning rate schedule. We report the results averaged over three independent runs, as shown in Figure 10.
Table 10: Cross-architecture results, rounded to one decimal place.
Model Nette Woof Fruit Yellow Meow Squawk
[…]
Mean 58.1 34.6 39.9 53.6 34.0 58.0
Table 11: Results on GSDD on CIFAR-10/100
@@
A.3 Parallelization Strategy for Batched Rendering
@@
Core Challenge and the Global ID System
@@
1D Flattening and Contiguous Memory Layout
@@
A.4 Visualization of Antialias And Gaussian Escape

### v2

@@ 4.3 Ablation Study
To investigate the contribution of each component to model performance, we conduct a series of ablation studies, with results summarized in Table 6. The experiments show that opacity modeling and bf16 mixed-precision training are two major factors, as removing either leads to a notable performance drop. The boundary entry in the table refers to the result obtained when both hard clipping and loss regularization are removed. The impact of anti-aliasing exhibits a dependence on the dataset and the complexity of the representation. Its effect becomes important when the number of Gaussians used to represent each image is limited. Specifically, for datasets like CIFAR-10, where storage budgets are constrained and each image is represented by a small number of Gaussians, disabling anti-aliasing introduces rendering artifacts that degrade performance. In contrast, for datasets like ImageNette with denser Gaussians, aliasing artifacts are largely mitigated, and the performance gains from anti-aliasing become relatively minor.
We further examine the effect of GPC through an ablation study, as shown in Figure 6. When using the memory-efficient DM algorithm or when distilling under small storage budgets, performance increases with GPC at first but eventually declines. This is because small GPC values lead to insufficient diversity, whereas excessively large GPC values allocate too few Gaussians per image, weakening each image’s representational capacity. In more memory-consuming settings, the optimal turning point of GPC becomes difficult to reach, and performance tends to improve monotonically as GPC increases. Figure 6 also shows that the performance of the initial dataset serves as a useful reference for estimating a reasonable GPC range.
Figure 6: Effect of GPC (Gaussian Images Per Class) on distilled dataset performance across different datasets, storage budgets, and distillation algorithms (TM Cazenavette_2022_CVPR and DM zhaoDatasetCondensationDistribution2023). Initial Accuracy denotes the performance of initialized gaussian images (initialized on sampled real images). Distilled Accuracy denotes the performance after distillation.
@@
Appendix A Experimental Details
A.1 Datasets
@@ A.4 Dataset Distillation
Table 7: Hyperparameters for dataset distillation on CIFAR-10 $32\times 32$.
@@
Table 8: Hyperparameters for dataset distillation on CIFAR-100 $32\times 32$.
@@
Table 9: Hyperparameters for dataset distillation on ImageNet-subset TM $128\times 128$.
@@
Table 10: Hyperparameters for dataset distillation on ImageNet-subset $128\times 128$ (DM).
@@
Table 11: Hyperparameters for dataset distillation on ImageNet-subset $256\times 256$ (DM).
Setting GPC num_points batch_real batch_syn zca Iteration
IPC=1 120 364 512 512 False 20000
Table 12: Hyperparameters for dataset distillation on ImageNet-subset $128\times 128$ (DC).
@@
A.5 Performance Benchmarking
In Table 5, we evaluate the image-fitting performance of FreD shinFrequencyDomainBasedDataset2023, SPEED weiSparseParameterizationEpitomic2023, DDiF shin2025distilling, and GSDD. We select 100 images from each class of the ImageNette dataset, resulting in a total of 1,000 images, all at a resolution of 128. Each method is constrained to the same storage budget, meaning that, on average, each image is represented using 963 floating-point parameters. For SPEED, we follow the parameter-counting formula provided in its original paper,
$\text{\#Params}=DK+1.5NHk+R(3D^{2}2+7D)+L(D+1),$
and, given the storage constraint, we solve for all feasible hyperparameter combinations. The final configuration used in our experiments is (D=32, N=32, k=135, R=2, L=64). We attempted to evaluate HaBa liuDatasetDistillationFactorization2022 as well, but its model structure cannot be solved inversely under the resolution of 128 and the storage budget of 963 parameters. FreD, DDiF, and GSDD exhibit clear advantages in terms of estimating and meeting the storage budget. For each method, we search over Adam and SGD optimizers and learning rates spaced exponentially by a factor of 10, and we report the highest PSNR achieved. All models are trained for 1,000 iterations using MSE loss. Each experiment is repeated three times with random initialization, and the average result is reported. The results demonstrate that our method achieves strong representational capability and computational efficiency.
Appendix B More Experimental Results
B.1 Performance Comparison on Low-Dimensional Datasets
@@
Table 13: Classification accuracy on CIFAR-10 and CIFAR-100 under different IPC settings.
@@
B.2 Detailed Cross-architecture Results
We evaluate the generalization ability of distilled data across different architectures, including AlexNet, VGG11, ResNet18, and ViT. The network implementations follow those provided in cazenavetteGeneralizingDatasetDistillation2023. For training, the learning rate is set to $1\times 10^{-3}$ for AlexNet, $5\times 10^{-5}$ for ViT, and $1\times 10^{-2}$ for the remaining architectures. All models are optimized using Adam with a cosine annealing learning rate schedule. We report the results averaged over three independent runs, as shown in Figure 14.
Table 14: Detailed cross-architecture performance comparison.
Test Net Method Nette Woof Fruit Yellow Meow Squawk
AlexNet TM 13.2 ± 0.6 10.0 ± 0.0 10.0 ± 0.0 11.0 ± 0.2 9.8 ± 0.0
IDC 17.4 ± 0.9 16.5 ± 0.7 17.9 ± 0.7 20.6 ± 0.9 16.8 ± 0.5 20.7 ± 1.0
FreD 35.7 ± 0.4 23.9 ± 0.7 15.8 ± 0.7 19.8 ± 1.2 14.4 ± 0.5 36.3 ± 0.3
DDiF 60.7 ± 2.3 36.4 ± 2.3 41.8 ± 0.6 56.2 ± 0.8 40.3 ± 1.9 60.5 ± 0.4
GSDD 63.1 ± 1.3 33.6 ± 0.9 41.9 ± 2.0 53.5 ± 0.9 38.1 ± 0.8 60.1 ± 0.2
VGG11 TM 17.4 ± 2.1 12.6 ± 1.8 11.8 ± 1.0 16.9 ± 1.1 13.8 ± 1.3
IDC 19.6 ± 1.5 16.0 ± 2.1 13.8 ± 1.3 16.8 ± 3.5 13.1 ± 2.0 19.1 ± 1.2
FreD 21.8 ± 2.9 17.1 1.7 12.6 ± 2.6 18.2 ± 1.1 13.2 ± 1.9 18.6 ± 2.3
DDiF 53.6 ± 1.5 29.9 ± 1.9 33.8 ± 1.9 44.2 ± 1.7 32.0 ± 1.8 37.9 ± 1.5
GSDD 56.8 ± 1.6 32.7 ± 1.4 34.1 ± 3.0 57.3 ± 4.5 31.3 ± 1.8 55.9 ± 3.4
ResNet18 TM 34.9 ± 2.3 20.7 ± 1.0 23.1 ± 1.5 43.4 ± 1.1 22.8 ± 2.2
IDC 43.6 ± 1.3 23.2 ± 0.8 32.9 ± 2.8 44.2 ± 3.5 28.2 ± 0.5 47.8 ± 1.9
FreD 48.8 ± 1.8 28.4 ± 0.6 34.0 ± 1.9 49.3 ± 1.1 29.0 ± 1.8 50.2 ± 0.8
DDiF 63.8 ± 1.8 37.5 ± 1.9 42.0 ± 1.9 55.9 ± 1.0 35.8 ± 1.8 62.6 ± 1.5
GSDD 59.1 ± 3.0 39.5 ± 2.0 42.6 ± 0.8 55.9 ± 1.2 40.9 ± 2.9 62.6 ± 0.6
ViT TM 22.6 ± 1.1 15.9 ± 0.4 23.3 ± 0.4 18.1 ± 1.3 18.6 ± 0.9
IDC 31.0 ± 0.6 22.4 ± 0.8 31.1 ± 0.8 30.3 ± 0.6 21.4 ± 0.7 32.2 ± 1.2
FreD 38.4 ± 0.7 25.4 ± 1.7 31.9 ± 1.4 37.6 ± 2.0 19.7 ± 0.8 44.4 ± 1.0
DDiF 59.0 ± 0.4 32.8 ± 0.8 39.4 ± 0.8 47.9 ± 0.6 27.0 ± 0.6 54.8 ± 1.1
GSDD 53.4 ± 0.6 32.6 ± 1.1 41.0 ± 1.2 47.6 ± 0.8 25.8 ± 0.4 53.5 ± 2.1
B.3 Different Initialization Strategy
Table 15: Initialization Strategy Ablation Results
Init Acc.
IDC real 61.4
FreD real 66.8
NSD random 68.6
DDiF real 72.0
GSDD random 63.6
GSDD warmup 69.4
GSDD real 76.4
Prior methods often initialize distilled data using real images, which may introduce additional privacy risks. We further conduct an ablation study on different initialization strategies, including random initialization. However, we find that optimizing Gaussian representations from purely random initialization is inherently unstable: the initial Gaussian points are extremely small and sparsely distributed, creating large gaps that prevent effective gradient propagation. To mitigate this issue, we introduce a Solid-Color Warmup strategy. Instead of fitting any real images, all Gaussian representations are first optimized to fit a single solid-color image. This process reduces point sparsity and adjusts the initial point sizes, after which standard distillation begins. As shown in Table 15, even without real-image initialization, our warmup-based strategy achieves strong results and surpasses several methods that depend on real-image initialization.
B.4 Results on ImageNet-1K
Table 16: Performance on ImageNet-1K.
IPC=1 IPC=10
SRe2L 0.9 21.3
SRe2L+IDC 13.4 41.5
SRe2L+GSDD 29.0 46.2
Table 17: Hyperparameters for each method on ImageNet-1K.
IPC=1 IPC=10
SRe2L+IDC k=20 k=5
SRe2L+GSDD GPC=50 GPC=200
Enabled by the efficiency of recent dataset-distillation frameworks, such as SRe2L yinSqueezeRecoverRelabel2023 and RDED sunDiversityRealismDistilled2024, dataset parameterization at the ImageNet-1K 5206848 scale has become feasible for the first time. In this work, we build upon SRe2L and evaluate our method under both IPC=1 and IPC=10 settings. In addition, we reproduce the IDC kimDatasetCondensationEfficient2022 baseline, where each image is spatially downscaled by a factor of $\sqrt{k}$, resulting in $k$-times more synthetic samples.
For IDC, under IPC=1 we report the best-performing $k$ selected from $k\in\{4,20,50\}$, and under IPC=10 from $k\in\{4,10,20\}$. As shown in Table 17, our method improves the performance of SRe2L on ImageNet-1K, particularly in the IPC=1 setting. Although the gains diminish at IPC=10, our approach still maintains a clear margin over IDC.
Appendix C Parallelization Strategy for Batched Rendering
@@
C.1 Core Challenge and the Global ID System
@@
C.2 1D Flattening and Contiguous Memory Layout
@@
C.3 Benchmark across different GPUs
To evaluate the stability of our CUDA-based rasterizer, we conduct systematic benchmarking on five different GPU architectures. We use an image-fitting task in which Gaussian representations reconstruct 1000 images sampled from ImageNet (10 classes $\times$ 100 images in total). Each experiment is repeated three times, and we average all results. We test four storage budgets (21, 170, 682, 2730 Gaussians per image). All experiments used PyTorch 2.5.1 and CUDA $\leq$ 12.0.
Table 18: Runtime, memory usage, and PSNR of our rasterizer benchmarked across five GPU architectures under varying storage budgets.
Time (s) Memory (GB) PSNR
GPU Model 21 170 682 2730 21 170 682 2730 21 170 682 2730
Tesla-V100-SXM2-32GB 23.65 26.11 30.28 42.94 1.66 1.70 1.81 2.28 20.24 25.93 30.65 36.09
NVIDIA-A100-SXM4-40GB 10.45 12.87 19.09 41.53 1.66 1.70 1.81 2.28 20.24 26.03 31.06 36.08
NVIDIA-GeForce-RTX-3090 10.33 16.07 25.44 61.38 1.66 1.70 1.81 2.28 20.24 26.03 31.06 36.08
NVIDIA-GeForce-RTX-4090 7.60 9.38 14.06 31.00 1.66 1.70 1.81 2.28 20.24 26.02 31.03 36.06
NVIDIA-GeForce-RTX-5090 7.13 6.99 9.54 21.79 1.66 1.70 1.81 2.28 20.23 26.02 31.03 36.06
As shown in Table 18. Different GPUs exhibit different runtimes due to heterogeneous compute capabilities, and RTX5090 is the fastest, while RTX3090 is the slowest. Memory usage remained identical across all platforms, and PSNR variation is small across architectures. Slight performance drops on V100 are expected due to older hardware, but results on A100/3090/4090/5090 remain highly consistent.
C.4 Results on Cross-Resolution Generalization
Gaussian Splatting naturally supports rendering at arbitrary resolutions through supersampling, enabling us to distill data at a low resolution (e.g., 128) while still using it to train models at a higher resolution (e.g., 256). This effectively reduces the computation when distilling datasets at large resolutions. Following results and setup in DDiF shin2025distilling, we conduct cross-resolution experiments as shown in Table 19.
Table 19: Test accuracies with different resolutions and networks. Images are first distilled with low resolution(128) and test at higher resolution(256/512).
test resolution test network method accuracy↑ difference↓ ratio↓
256 ConvNetD5 Vanilla 31.2 20.2 0.39
IDC 55.0 6.4 0.10
SPEED 58.8 8.1 0.12
FreD 56.4 10.4 0.16
DDiF 66.3 5.7 0.08
GSDD 66.4 10.0 0.13
ConvNetD6 Vanilla 44.0 7.3 0.14
IDC 55.4 6.0 0.10
SPEED 62.6 4.3 0.06
FreD 61.8 5.0 0.07
DDiF 70.6 1.4 0.02
GSDD 71.6 4.8 0.06
512 ConvNetD5 Vanilla 27.4 24.0 0.47
IDC 39.5 21.9 0.36
SPEED 45.0 21.9 0.33
FreD 42.9 23.9 0.36
DDiF 58.7 13.3 0.18
GSDD 59.4 17.0 0.22
ConvNetD6 Vanilla 41.2 10.1 0.20
IDC 51.5 9.9 0.16
SPEED 60.1 6.8 0.10
FreD 56.3 10.5 0.16
DDiF 69.0 3.0 0.04
GSDD 67.8 8.6 0.11
For IDC, SPEED, and FreD, we adopt their best-performing upsampling strategies. For DDiF and GSDD, supersampling is applied, allowing lossless upscaling of the distilled images. GSDD maintains competitive performance in cross-resolution tasks. Although its relative performance drop is larger than DDiF, which is likely due to GSDD’s inherent sparsity, reducing its ability to capture fine-grained features when evaluated at higher resolutions.
Appendix D Visualization of Antialias And Gaussian Escape